# Leveraging Consistent Spatio-Temporal Correspondence for Robust Visual Odometry

### v1

@@ Introduction
was proposed to combine the strengths of classical methods and deep learning methods. There has already been a series of outstanding works in this field such as (Teed and Deng 2021; Teed, Lipson, and Deng 2024; Chen et al. 2024; Gurumurthy et al. 2024; Klenk et al. 2024; Cheng et al. 2024a), among them, two particularly notable and closely related to our work are DROID-SLAM and DPVO. DROID-SLAM integrates RAFT for iterative dense optical flow prediction to obtain a more robust correspondence, which greatly improves the robustness of VO system. To improve the efficiency of VO, DPVO (Teed, Lipson, and Deng 2024) replaces dense optical flow tracking with patch-based tracking. However, this patch-based matching approach can exacerbate instability issues in extreme scenarios, such as motion blur or highly dynamic scenes.
@@
To leverage the benefits of temporal ordering, allowing the system to effectively utilize the sequential information and enhance temporal consistency, STVO additionally maintains a motion state ${m}^{k}_{t}\in{R}^{H\times W\times{D}_{m}}$ for each Source Edge Set $E_{t}$, where ${m}^{0}_{t}$ is randomly initialized. The motion state ${m}^{k}_{t}$ contains motion information at the $k$-th iteration for $I_{t}$ and will be updated in each iteration.
@@
In Table 4, we benchmark on the EuRoC MAV (Burri et al. 2016) dataset. EuRoC MAV dataset is a widely used benchmark for evaluating VO systems, collected using Micro Aerial Vehicles (MAV). Like DROID-VO and DPVO, we skip every other frame, doubling the effective frame rate of the system. STVO outperforms previous methods on the majority of sequences on EuRoC MAV. The average error of STVO is $5\%$ lower than the state-of-the-art DPVO method and $21\%$ lower than DROID-VO.

### v4

@@ Introduction
was proposed to combine the strengths of classical methods and deep learning methods. There has already been a series of outstanding works in this field such as (Teed and Deng 2021; Teed, Lipson, and Deng 2024; Chen et al. 2024; Gurumurthy et al. 2024; Cheng et al. 2024a; Klenk et al. 2024), among them, two particularly notable and closely related to our work are DROID-SLAM and DPVO. DROID-SLAM integrates RAFT for iterative dense optical flow prediction to obtain a more robust correspondence, which greatly improves the robustness of VO system. To improve the efficiency of VO, DPVO (Teed, Lipson, and Deng 2024) replaces dense optical flow tracking with patch-based tracking. However, this patch-based matching approach can exacerbate instability issues in extreme scenarios, such as motion blur or highly dynamic scenes.
@@
To leverage the benefits of temporal ordering, allowing the system to utilize the sequential information and enhance temporal consistency, STVO maintains a motion state ${m}^{k}_{t}\in{R}^{H\times W\times{D}_{m}}$ for each Source Edge Set $E_{t}$, where ${m}^{0}_{t}$ is randomly initialized. The motion state ${m}^{k}_{t}$ contains motion information at the $k$-th iteration for $I_{t}$ and will be updated in each iteration.
@@
In Table 4, we present our benchmarking results on the EuRoC MAV(Burri et al. 2016) dataset. EuRoC MAV is a widely used benchmark for evaluating VO systems, collected using Micro Aerial Vehicles (MAV). The dataset includes fast and dynamic camera movements that pose significant challenges for VO systems. To align with the practices of similar studies such as DROID-VO and DPVO, we process every other frame, effectively doubling the system’s frame rate to 40 FPS. STVO outperforms previous methods on the majority of sequences on EuRoC MAV. The average error of STVO is $5\%$ lower than the state-of-the-art DPVO method and $21\%$ lower than DROID-VO.